OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-7,-9) and (-1,-8)

OpenStudy (anonymous):

slope = [y(2)-y(1)]/[x(2)-x(1)] find the slope first. post it here and i'll give you the next part.

OpenStudy (anonymous):

1x-1y?

OpenStudy (anonymous):

for our x and y values we are going to use the points from above, lets make (-7,9) point number 1 and (-1,-8) point number two. so our points are set up like this: point 1 = (x(1),y(1)) = (-7,-9) point 2 = (x(2),y(2)) = (-1,-8) so using that data, and plugging it into the equation above, we get: [(-8)-(-9)]/[(-1)-(-7)] when we solve that out it will give us: -8+9/(-1+7) = 1/6 so the slope is 1/6. normally the slope won't have any variables in it.

OpenStudy (anonymous):

next we want to find the y-intercept. since we need that to put into our equation for the line. the equation for the line is y=mx+b. where m is the slope and b is the y-intercept. since we already know the slope, lets put that in: y=(1/6)x+b

OpenStudy (anonymous):

now we want to find the y-intercept. since we know two points that are on the line, we can just use one of those to find the y-intercept. lets use point 1 which was (-7,-9). since we know that the -7 is the x and the -9 is the y, we will put those into the equation we just wrote down: -9=(1/6)(-8)+b now we solve for b: -9=-(8/6)+b -9+(8/6)=b -9+(4/3)=b --reduce the fraction -(27/3)+(4/3)=b -- multiply the -9 by (3/3) so that we can combine the numbers -(23/3)=b now that we know b, we can put our complete equation together: y=(1/6)x-(23/3)

OpenStudy (anonymous):

now use the same method to work another problem!